O Assignment Score 1167/1700 9 Hint Resources Question 10 of 17> A speedboat moves on a lake with initial velocity vector V1 = 8.81 m/s and ul y 2.97 m/s, then accelerates for 6.71 s at an average acceleration of aw,x-0.107 m/s? and awvy 0.103 m/s2. What are the components of the speedboats final velocity, u2x and vzy? m/s m/s Find the speedboats final speed. final speed: m/s
Answer #1

By using, v = u + at

v2, x = v1, x + ax*t

= 8.81 + (-0.107)*6.71

= 8.1 m/s

v2, y = v1, y + ay*t

= - 2.97 + (0.103)*6.71

= - 2.28 m/s

Final speed = sqrt(8.1^2 + (-2.28)^2)

= 8.41 m/s